OpenStudy (anonymous):

hey, i've been stuck on this for like 2 weeks and i need some help understanding.

OpenStudy (anonymous):

robbie the robot is on a weather satellite orbiting earth about 3600km above the surface. the earth's radius is bout 6400km. he has had a malfunction in his output device, and the satellite is traveling without communication. his last report was only in terms of trigonometric values and was only partially received. it said sin theta<0 and then he was lost again. part 1: create a set of coordinates that would be reasonable for robbie's position in space and satisfy his last, partial report. using complete sentences, describe robbie's location and your reasoning

OpenStudy (anonymous):

it took me a few days, but i realized that you would add the radius and his orbit above earths surface, which is 6400+3600=10000 km and that he is on the circumference. i know sin is negative in quadrants 3 and 4 and that one of the coordinates would be negative but i am confused on what to do next in order to find the coordinates to use

OpenStudy (anonymous):

@ganeshie8 @Directrix

OpenStudy (anonymous):

@TheSmartOne @Kainui

OpenStudy (kainui):

I think you've done a good job so far and this problem is considerably open ended. I guess what I would do is something like this  So looking here and knowing that, since he's travelling at what I'm assuming to be a constant speed, maybe he's something like \[\Large x=10000\cos(180+ \omega t) \\ \Large y = 10000\sin(180+\omega t)\] Since any time greater than t=0 will have a negative value for sine, and it lets us predict his path too. Omega is the angular velocity, and t is time, so rate times time is distance, but it will be an angle theta since it's angular velocity. Does this seem like what you're looking for?

OpenStudy (anonymous):

kinda.. im just a little slow sometimes, can you explain it a little simpler please? so the t is time but what does the w stand for?

OpenStudy (kainui):

That w is a greek letter omega. It's just the speed. By analogy, here we have distance is rate times time and there we have angle is angle speed times time. \[\Large d= r*t \\ \Large \theta = \omega * t\] So if you substitute this into those equations there we have sin(180+ theta) which just means that we have the 180 degrees to make sine negative and then theta is the extra distance it went after we lost contact with it. I don't mean to be confusing I just sort of use some of these technical words too often that I forget not everyone knows them, so don't feel bad. =P

OpenStudy (anonymous):

lol its okay :). and okay, so now i understand that part but i have another question. what would that mean in terms of finding the coordinates? is there a way to figure out hat theta actually is in this situation?

OpenStudy (anonymous):

@waterineyes @confluxepic

OpenStudy (kainui):

Well I mean the only other option would be to guess that it started falling towards the earth instead of remaining in orbit, but that would get more complicated than I think the question is asking for. Above I listed x= and y= to say those are the coordinates relative to the center of the earth if that's what you are looking for. Other than that, I really don't have anything more to say.

OpenStudy (anonymous):

oh, okay, thanks :)